How many grams of water are produced when a 4 g sample of methane CH_4 is burned according to the balanced equation?
CH_4 + 2 O_2 → CO_2 + 2 H_2O
Answer #1

Molar mass of CH4,

MM = 1*MM(C) + 4*MM(H)

= 1*12.01 + 4*1.008

= 16.042 g/mol

mass of CH4 = 4 g

mol of CH4 = (mass)/(molar mass)

= 4/16.04

= 0.2493 mol

According to balanced equation

mol of H2O formed = (2/1)* moles of CH4

= (2/1)*0.2493

= 0.4987 mol

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

mass of H2O = number of mol * molar mass

= 0.4987*18.02

= 8.984 g

Answer: 8.98 g